OpenStudy (ledah):

Simplify the following ADDITION of radicals expression. sqrt25x+sqrt25y+sqrt16x

OpenStudy (ledah):

that doesn't tell me anything about how to work this.

OpenStudy (ledah):

oh.

OpenStudy (ledah):

well I don't think that's what it's asking. because 14 isn't an answer in this problem

OpenStudy (ledah):

its asking to simplify the addition of the radicals

OpenStudy (anonymous):

9x + 5y

OpenStudy (anonymous):

or 9√x + 5√y

OpenStudy (anonymous):

@Ledah Your post's confusing!

OpenStudy (anonymous):

Yup. \[\sqrt{25x} + \sqrt{25y} + \sqrt{16x} = 5\sqrt{x} + 5\sqrt{y} + 4\sqrt{x} = 9\sqrt{x} + 5\sqrt{y}\]

OpenStudy (ledah):

I thought it was pretty clear. but then again. I'm probably the only one who understood it. so yeah I get where your coming from.

OpenStudy (anonymous):

The only thing that's confusing is whether the variables are under the radicals or not.

OpenStudy (anonymous):

I assume √25x + √16x + √25y

OpenStudy (ledah):

Yeah that would be it @Chlorophyll

OpenStudy (anonymous):

There you go! I'm glad at least you understand problem with your post!

OpenStudy (ledah):

yeah. sorry about that.

OpenStudy (anonymous):

Then the answer is: 9√x + 5√y